# Accurate Low Rank Approximation at a Low Computational Cost

### v13

@@ 1.3 LRA by means of escalation
[15, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
@@
1.3.2. Gaps in the spectrum and escalation of LRAs. For a fixed $r$ in a target range, we fix an upper rank $\rho$ satisfying $r\leq\rho\ll n$ and then (i) compute a rank-$\rho$ approximation of $M$ given by a triplet $\{X,Y,Z\}$ such that
$||M-M(\rho)||=\beta\sigma_{\rho+1}(M),~{}{\rm for}~{}M(\rho)=XYZ,~{}1\leq\beta\leq\beta_{+}~{}{\rm and~{}a~{}fixed~{}bound}~{}\beta_{+},$ (1.3)
and (ii) compute and output the $r$-top SVD of $M(\rho)$.
The relative error random sampling algorithms of [20, 54] use from $O(mn\log(\rho))$ to $O(mn\rho)$ flops and about $(m+n)\rho$ scalars to perform stage (i), while their heuristic variations in our Sec. 4 perform stage (i) at a lower cost; deterministic Alg. B.1 of Appendix B uses $O(m\rho^{2})$ flops and about $(m+n)\rho$ scalars to perform stage (ii). [^7] By applying the algorithm of [14] instead of Alg. B.1 at stage (ii), one can a little accelerate the computations while only slightly increasing the error norm bound.
In Alg. 2.1 of Sec. 2 we elaborate upon this outline and deduce that
$||M-(M(\rho))_{r}||\leq\sigma_{r+1}(M)+2\beta\sigma_{\rho+1}(M)=\Big{(}1+\frac{2\beta\sigma_{\rho+1}(M)}{\sigma_{r+1}(M)}\Big{)}\sigma_{r+1}(M).$ (1.4)
The output error norm $||M-M(\rho)_{r}||$ of Alg. 2.1 is small if the value $\frac{\sigma_{r+1}(M)}{2\beta\sigma_{\rho+1}(M)}$ is large, that is, if there is a large gap in the segment $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of the spectrum of $M$.
For comparison, a rank-$r$ approximation of $M$ is only interesting if the value $\frac{\sigma_{1}(M)}{\sigma_{r+1}(M)}$ is large, that is, if there is a gap in the leading segment $[\sigma_{1}(M),\sigma_{r+1}(M)]$ of the spectrum.
We can reveal gaps in the spectrum by approximating the values of $||M-(M(\rho))_{r}||$ for various pairs of $\rho$ and $r$ and comparing the output error norms.
1.3.3. Superfast variations. We can compute the $r$-truncation of $M(\rho)$ superfast for any matrix $M$ if $\rho$ is small such that $\rho^{2}=o(n)$, while some superfast variations of the known LRA algorithms (e.g., those of [20, 54] discussed in our in Sec. 4) output reasonably accurate rank-$\rho$ approximations $M(\rho)$ for a large class of FDS matrices $M$.
We cannot estimate the spectral error norm of a matrix superfast, and even its LAPACK heuristic estimation is not superfast, but we propose a heuristic superfast 1-norm estimator in Sec. 5 and test it in Sec. 6.4.
1.3.4. Iterative LRAs. In Alg. 3.1 of Sec. 3, which is a competitive alternative to Alg. 2.1, we write $X_{0}:=0$, $E_{i}:=M-X_{i}$, and $X_{i}:=(X_{i-1}+E_{i}(\rho))_{r}$ for $i=1,2,\dots,k$. We compress the matrices $X_{i-1}+E_{i}(\rho)$, that is, keep their ranks bounded by $r$ – by working with their $r$-truncation, and we do this superfast as long as their rank bound $\rho+r$ is low. For computation of the matrix $E_{i}(\rho)$ we use the algorithms of stage (i) of Alg. 2.1. This way we extend to LRA the popular iterative refinement algorithms.
The algorithm is essentially reduced to $k$ applications of Alg. 2.1.
In our tests Alg. 3.1 output LRAs with relative output errors about 1.1 for a matrix class marked Shaw for which Alg. 2.1 failed, but for all other inputs Alg. 3.1 was less accurate than Alg. 2.1 applied with the upper rank $\rho$ twice as large versus Alg. 2.1.
Our upper estimates for the accuracy of both algorithms are close to one anaother but are both pessimistic, and it seems that further comparative study of the two algorithms is in order.
For completeness, we cover iterative refinement of LRA with no compression in Appendix F, but unlike Alg. 3.1, we cannot provide any support for its numerical stability.
1.4 Related works
Motivated by exposition of the progress in LRA in [20, 54], we extend the rudimentary study of LRA escalation in [40, 41, 43, 42] (cf. the end of Sec. 4).
The algorithms in [25, 39] extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [51, Secs. 3.3.4 and 4.2.5], [19, Ch. 12 and Sec. 20.5], [7], [15, Secs. 3.5.3 and 5.3.8], [3, Secs. 1.4.6, 2.1.4, and 2.3.7]) but also widely applied in various other linear and nonlinear matrix computations (see [51, page 223 and the references on page 225]). Its extension to LRA is not straightforward – because the rank of the computed approximations grows recursively; we block this growth by applying re-compression, tracing such techniques to [31, 32, 37, 46, 38, 35] and [34, Secs. 6.3-6.5].
Very different iterative refinement of LRA in [25] uses no crude auxiliary LRA but instead bounds the Principal Angle Distances between singular spaces of $M$ and its LRAs and incorporates randomization based on the leverage scores of [8].
Historically, we first succeeded with iterative refinement of LRA and then noticed that escalation alone is quite competitive.
In our choice of abridged SRHT matrices for sparse multipliers we follow [40, 41, 42, 43], which proved some crude but reasonable probabilistic upper bounds on output errors of some superfast LRA algorithms under randomization in the input matrix space defined by a fixed probability structure in that space. [^8] These papers continued the study in [45, 44, 47], which introduced a probability structure in the space of input matrices to motivate application of Gaussian elimination without pivoting.
The relevance of the gaps in the spectrum of the singular values of a matrix to its LRA is well-known but has not been exploited in the way that we propose, test, and analyze.
Likewise, from the comments of an expert on an earlier draft of this paper, we learned that [9] also worked with error matrices of LRA, but in the directions quite different from ours.
1.5 Organization of our paper
We present our LRA escalators in Secs. 2 and 3 and discuss superfast computation of a crude LRA in Sec. 4. We specify a superfast heuristic 1-norm estimator in Sec. 5, cover our tests in Sec. 6, and devote our brief Sec. 7 to conclusions. In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails. In Appendix B we compute $r$-top SVD of an LRA or any low-rank matrix superfast. In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA. In Appendix D we generate abridged SRHT matrices. In Appendix E we briefly recall ACA (C-A) iterations. In Appendix F we iteratively refine LRA with no compression.
2 LRA by means of escalation
Next we elaborate upon escalation outlined in Sec. 1.3.2.
@@ 2 LRA by means of escalation
Two positive integers $m$ and $n\leq m$, an $m\times n$ matrix $M$, possibly given implicitly by its sketches, three positive values $\widehat{\tau}$, $\widehat{\mu}$, and $r$, $r\ll n$, denoting bound on the overall computational cost, [^9] Here and hereafter we use a single character $\tau$ to represent both flops and memory size involved. an absolute error norm bound, and a fixed integer in a target range, respectively; three auxiliary Algs. $\mathbb{A}$, $\mathbb{B}$, and $\mathbb{C}$ which for any $\rho$, $r\leq\rho\leq n$, compute a (crude but reasonably close) rank-$\rho$ approximation of $M$, compute $r$-top SVD of a rank-$\rho$ matrix of size $m\times n$, and estimate the norm of the difference of $M$ and another $m\times n$ matrix, respectively, by performing within computational cost bounds $\tau(\rho)$, $\tau(\rho,r)$, and $\tau^{\prime}$, respectively. [We elaborate upon Algs. ${\bf\mathbb{A}}$, ${\bf\mathbb{C}}$, and ${\bf\mathbb{B}}$ in Secs. 4 and 5 and Appendix B, respectively.]
@@
FAILURE or SVD of a rank-$r$ matrix $X\approx M$ and the norm $||M-X||$ not exceeding $\widehat{\mu}$.
@@
Fix the maximal upper rank $\rho$ satisfying the following bounds (see Remark 2.1),
$\tau(\rho)+\tau(\rho,r)+\tau^{\prime}\leq\widehat{\tau}~{}{\rm and}~{}r\leq\rho\ll n\leq m.$ (2.1)
@@
    Apply Alg. $\mathbb{A}$ to output a rank-$\rho$ approximation of $M$ given by SVD of some matrix $M(\rho)$.
@@
    Apply Alg. ${\bf\mathbb{B}}$ to that matrix to compute its $r$-top SVD,101010 Alg. B.1 of Appendix B performs this stage superfast where $\rho^{2}\ll n$. which is SVD of $X:=M(\rho)_{r}$.
@@
    If Alg. ${\bf\mathbb{C}}$ shows that
    $\mu:=||E||\leq\widehat{\mu}~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
    then output that SVD and the estimated norm $\mu$. Otherwise output FAILURE.
Remark 2.1.
[The computational cost bound, the choice of an upper rank, and a simplified version of Alg. 2.1.] (i) Bound (2.1) implies that the overall computational cost $\tau(\rho)+\tau(\rho,r)+\tau^{\prime}$ stays within tolerance $\widehat{\tau}$. (ii) Only limited by Eqn. (2.1), we maximize the upper rank $\rho$ – to decrease the error norm bound $\mu$. (iii) By increasing the upper bound $\widehat{\mu}$ on the output error norm $\mu$, we can decrease the cost of computations at stage 2, which, however, tends to be dominated. (iv) In Sec. 6.3 we test a simplified version of Alg. 2.1 where we impose no upper bounds $\widehat{\tau}$ and $\widehat{\mu}$ and display the output error norms for $\rho:=ir$ and $i=2,3,4,5$.
@@
For $\mu$ of Eqn. (2.2) and $\beta$ of (1.3), it holds that
$\mu\leq\sigma_{r+1}(M)+\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||=\beta\sigma_{\rho+1}(M).$ (2.3)
@@
Represent $E$ of (2.2) as $M-M_{r}+M_{r}-M(\rho)_{r}$, apply triangle inequality, and obtain
$||E||\leq||M-M_{r}||+||M_{r}-M(\rho)_{r}||=\sigma_{r+1}(M)+||M_{r}-M(\rho)_{r}||.$
Thm. 1.2 implies that $||M_{r}-M(\rho))_{r}||\leq\nu$. Combine the latter two bounds. ∎
Corollary 2.1.
[Relative output error norms of Algs. $\mathbb{A}$ and 2.1.] Write (cf. (1.3))
$\alpha:=\alpha(M,\rho,r):=\frac{||E||}{\sigma_{r+1}(M)}-1~{}{\rm and}~{}\beta:=\beta(M,\rho):=\frac{\nu}{\sigma_{\rho+1}(M)}.$ (2.4)
Then the rank-$r$ approximation of $M$ output by Alg. 2.1, for a fixed pair of $r$ and $\rho$, is optimal up to within the absolute and relative error norms $\nu$ and $1+\alpha$, respectively, such that
$1+\alpha\leq\frac{\nu}{\sigma_{r+1}(M)}=\frac{\beta\sigma_{\rho+1}(M)}{\sigma_{r+1}(M)}.$ (2.5)
@@
How sharp are the upper bounds (2.3)-(2.5), that is, how much can we lose when we apply triangle inequality and Thm. 1.2 in the proof of Thm. 2.1? The loss can be significant: substitute $\rho=r$ into (2.3) and obtain $\mu:=||M-M(\rho)_{r}||\leq(1+\beta)\sigma_{r+1}(M)$, but more directly obtain $\mu:=||M-M(\rho)_{r}||=||M-M(r)||=\beta\sigma_{r+1}(M)$ because $M(\rho)_{r}=M(r)$.
3 Iterative escalation with compression
Next we elaborate upon iterative escalation with compression, sketched in Sec. 1.3.4.
@@ 3 Iterative escalation with compression
as in Alg. 2.1, except that the input of Alg. 3.1 additionally includes a positive integer $k$ and an Alg. $\mathbb{S}$ that adds or subtracts a pair of black box $m\times n$ matrices at a computational cost $\tau(\pm)$.
@@
    Apply Alg. ${\bf\mathbb{A}}$ to the matrix $E_{i}$ and the upper rank $\rho$. [The algorithm outputs a rank-$\rho$ approximation of $E_{i}$ given by SVD of a matrix $E_{i}(\rho)$.]
@@
    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$.
@@
    Apply Alg. ${\bf\mathbb{B}}$ to this matrix and the fixed $r$ to compute $r$-top SVD of $M_{i}(\rho)$, that is, SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$.
@@
    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if Alg. ${\bf\mathbb{C}}$ shows that $2\mu_{i}:=2\mu_{i}(M,\rho)\leq\widehat{\mu}$.
    Otherwise output FAILURE if $i=k-1$.
@@
    Otherwise compute $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1.
@@
[The computational cost bound, the choice of an upper rank, and a simplified version of Alg. 3.1.] (i) (3.1) implies that the overall computational cost $\tau(\rho)+\tau(\rho,r)+\tau^{\prime}$ stays within tolerance $\widehat{\tau}$. (ii) As in Alg. 2.1 we maximize the upper rank $\rho$ to decrease the output error norm, being only limited by a bound on the overall computational cost. (iii) In principle, we can allow an upper rank $\rho$ to vary as $i$ varies; then we can readily extend the analysis of the algorithm. (iv) In Sec. 6.3 we test a simplified version of Alg. 3.1 where we impose no upper bounds $\widehat{\tau}$ and $\widehat{\mu}$, fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ elsewhere.
In the following extension of Thm. 2.1 we estimate the output error norm of Alg. 3.1.
@@
For the error matrices $E_{i+1}$ of Alg. 3.1, $\nu_{i}:=||E_{i}-E_{i}(\rho)||$, $i=0,1,\dots,k-1$, and an upper bound
$\beta\geq\beta_{i}(M,\rho):=\frac{\nu_{i}}{\sigma_{\rho+1}(E_{i})}~{}{\rm for}~{}i=0,1,\dots,k-1$ (3.2)
on the relative errors of Alg. $\mathbb{A}$ (unlike (1.3) we write $\beta$ rather than $\beta_{+}$), it holds that
$\mu_{i+1}:=||E_{i+1}||\leq||M-M_{r}||+\nu_{i}=\sigma_{r+1}(M)+\beta\sigma_{\rho+1}(E_{i}).$
@@
Recall from Alg. 3.1 that $E_{i+1}=M-X_{i+1}$ for $X_{i+1}=(X_{i}+E_{i}(\rho))_{r}$. Hence $E_{i+1}=M-M_{r}+M_{r}-X_{i+1}$, and so $||E_{i+1}||\leq||M-M_{r}||+||M_{r}-X_{i+1}||=\sigma_{r+1}(M)+||M_{r}-X_{i+1}||$.
It remains to prove that
$||M_{r}-X_{i+1}||\leq\beta\sigma_{\rho+1}(E_{i}).$ (3.3)
Since $X_{i+1}=(X_{i}+E_{i}(\rho))_{r}$, it follows that $M_{r}-X_{i+1}=M_{r}-(X_{i}+E_{i}(\rho))_{r}$. Combine this equation with Thm. 1.2 and obtain that $||M_{r}-X_{i+1}||\leq||M-(X_{i}+E_{i}(\rho))||=||E_{i}-E_{i}(\rho)||$. Now bound (3.3) and hence the theorem follow from Eqn. (3.2). ∎
@@
Extend Cor. 2.1 by writing $\alpha_{i+1}$ and $E_{i+1}$ instead of $\alpha$ and $E$. Then
$\alpha_{i+1}:=\alpha_{i+1}(M,r,\rho):=\frac{||E_{i+1}||}{\sigma_{r+1}(M)}-1\leq\beta\frac{\sigma_{\rho+1}(E_{i})}{\sigma_{r+1}(M)},~{}i=0,1,\dots,k-1.$ (3.4)
Lemma 3.1.
Recall that $\sigma_{j}(X)=0$ for $j>n$ and write $\sigma_{j}(X)=\infty$ for any $m\times n$ matrix $X$ and $m\geq n$. Then $\sigma_{j+r+1}(M)\leq\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$ for $E_{i}$ of Alg. 3.1 and all $i$ and $j$.
Proof.
Recall that $\operatorname{rank}(M-X_{i})\leq r$ for all $i\geq 0$, $\sigma_{j+1}(E_{i})$ is an optimal rank-$j$ approximation of the matrix $E_{i}$ by virtue of Thm. 1.1, and rank-$\rho$ approximation of $M$ with $Y$ does not exceed rank-$(\rho-r)$ approximation of $M$ with $Y+X$ if $\operatorname{rank}(X)=r$. Combine these observations. ∎
Combine bound (3.4) with the lemma for $j=\rho+1$ to obtain
Corollary 3.2.
The singular values $\sigma_{\rho+1}(E_{i})$ of Eqn. (3.4) satisfy
@@
[Error norm bounds.] For $E_{i+1}=E$ and $E_{i}=M$, bound (3.4) turns into bound (2.5) of Cor. 2.1, and this is precisely what we have at the initial stage for $i=0$. For $i>0$ bound (3.4) can be weaker or stronger than that, but it cannot be stronger than (2.5) with $\rho$ replaced by $\rho+r$. In our tests for Algs. 2.1 and 3.1 in Sec. 6.3 for for matrices Shaw and $r=20$ Alg. 3.1 for $\rho=2r$ superseded Alg. 2.1 for $\rho=ir$, $i=3,4,5$. This implies that the upper bound of Cor. 3.2 is overly pessimistic for worst case inputs, similarly to bound of Cor. 2.1 (cf. Remark 2.2).
4 LRA by means of random sampling with sparse multipliers
Alg. 1.1 is near-optimal with a high probability (whp) but not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT [20], [54, Sec. 3.9].
One can accelerate it to a superfast level by using Ultra-Sparse Rademacher test matrices [54, Sec. 3.9]), [6], which depend on an integer parameter $s$, defining sparsity. They are sparse and support superfast LRAs for smaller values $s$.
Likewise, in Appendix D we turn SRHT test matrices into abridged SRHT test matrices of a small length, which support superfast LRA; this recipe can be extended to various other recursively generated test matrices such as SRFT matrices and those of [20, Remark 4.6], [15, Sec. 5.1.8].
According to [22], such acceleration tends to make the output LRAs somewhat less reliable. Apparently for this reason, our tests for synthetic inputs from [55, Sec. 7.3.1] failed with abridged SRHT test matrices of length 3, while the same tests succeeded with Gaussian test matrices. For the other inputs, however, our tests succeeded about equally well with test matrices of both kinds, even where the output LRAs of Alg. $\mathbb{A}$ were much cruder with abridged SRHT test matrices than with Gaussian random ones.
[40, 41, 42, 43] studied primitive algorithms for CUR LRAs (cf. Appendix C.1) generated by pairs of test matrices made up of two random subsets of $k$ rows and $\ell$ columns of the identity matrix. These algorithms run at very low cost and output rather crude LRAs, and so [40, 41, 42, 43] proposed to strengthen them with escalation, by computing the sets of $k$ rows and $\ell$ columns as proper subsets of the sets of $p>k$ random rows and $q>\ell$ random columns of the identity matrix. [40, 41, 42, 43] call these algorithms cynical – representing a minor advance from the primitive ones. Not competitive for LRA, such cynical algorithms were predecessors of our use of escalation and could still be considered for Alg. $\mathbb{A}$.
5 Superfast 1-norm estimator
Next we propose a superfast variation of the LAPACK estimator for 1-norm $||E||_{1}$ of a matrix $E\in\mathbb{R}^{m\times n}$ (cf. [19, Alg. 15.3], [51, Alg. 5.3.1]) and recall that $||E||_{1}=||E^{T}||_{\infty}$, $||E||_{\infty}\leq n||E||$, $||E||_{1}\leq m||E||$, and $||E||\leq||E||_{1}||E||_{\infty}$ (see [15, Eqns. (2.3.11), (2.3.12) and Cor. 2.3.2])).
Empirically, the LAPACKl norm-1 estimator converges in about 4.5 iterations on the average and is superfast at every iteration, except for its initial and verification stages, at which $E$ is multiplied by dense auxiliary vectors. We accelerate these stages by using sparse vectors instead. We keep a small number of randomly sampled components of the original vectors, set the remaining components to 0, and scale the resulting vectors to turn their 1-norm into 1; we only test our 1-norm estimator for matrices $E$ representing the errors of LRA.
Algorithm 5.1.
[A superfast 1-Norm Estimator.]
Input:
An $m\times n$ matrix $E$, a positive integer $k\ll n\leq m$, and a positive tolerance TOL.
Output:
A vector ${\bf u}$ such that $||E{\bf u}||_{1}\approx||E||_{1}$.
Initialization:
Generate a sparse unit vector ${\bf v}$ such that $||{\bf v}||_{1}=1$ as follows: choose $k$ random integers $i_{1},\dots,i_{k}$ in the range from 1 to $n$ and let ${\bf v}=\frac{1}{k}(v_{i})_{i=1}^{n}$ where $v_{i_{j}}=1$ for $j=1,\dots,k$ and $v_{i}=0$ for all other $i$. Write COUNTER:=0.
Computations:
  1. 1.
    Compute the vectors ${\bf u}=E{\bf v}$, ${\bf w}=(\operatorname{sign}(u_{i}))_{i=1}^{n}$, and ${\bf x}=E^{T}{\bf w}$.
  2. 2.
    If $||{\bf x}||_{\infty}\leq||{\bf u}||_{1}$, output ${\bf u}$ and terminate.
  3. 3.
    Choose $j$ such that $|x_{j}|=||{\bf x}||_{\infty}$ and let ${\bf v}={\bf e}_{j}$. Increase COUNTER by 1. Go to stage 1 if COUNTER$<$Tol. Otherwise stop and output FAILURE.
We also apply the algorithm to a small number of other choices of the initial sparse unit vector ${\bf v}$ and output ${\bf u}$ that maximizes the norm $||E{\bf u}||_{1}$. For these choices we first define vectors of dimension $n$ sharing $k$ components on $k$ random positions with the following vector
$\left(1,-1-\frac{1}{n-1},1+\frac{2}{n-1},1-\frac{3}{n-1},\dots,(-1)^{n-1}2\right)^{T}$
and otherwise filled with 0s; then scale them to make their 1-norm equal to 1.

### v16

@@ 2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
[7, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
@@
In the following algorithm we elaborate upon the outline of Sec. 1.2 assuming that we are given two black box subroutines that output a crude LRA of a matrix.
@@
An $m\times n$ matrix $M$, for $n\leq m$, given implicitly by an oracle (black box subroutine) for its multiplication by a vector (cf. Sec. 1.1.1), a target rank $r\ll n$, and a tolerance $\mu$ to the output error norm.
@@
FAILURE or SVD of a rank-$r$ matrix $X\approx M$ such that $||M-X||\leq\mu$.
@@
Fix an upper rank $\rho$ satisfying
$r\leq\rho\ll\sqrt{n}.$ (2.1)
@@
    Compute an approximation of $M$ given by SVD of a matrix $M(\rho)$ of rank $\rho$.
@@
    Compute its $r$-top SVD (cf. Appendix B), which is SVD of $X:=M(\rho)_{r}$.
@@
    Estimate the norm $||E||$. If
    $||E||\leq\mu~{}{\rm for}~{}E:=M-X=M-M(\rho)_{r},$ (2.2)
    then output that SVD and $||E||$. Otherwise output FAILURE.
@@
The error norm $||E||=||M-X||$ (cf. Eqn. (2.2)) satisfies
$||E||\leq\sigma_{r+1}(M)+2\nu~{}{\rm where}~{}\nu:=\nu(M,\rho):=||M-M(\rho)||,$ (2.3)
which turns into (1.3) for $\sigma_{r+1}(M)=||M-M_{r}||$.
@@
Represent $E$ of (2.2) as $M-M(\rho)+M(\rho)-M(\rho)_{r}$, apply triangle inequality, and obtain
$||E||\leq||M-M(\rho)||+||M(\rho)-M(\rho)_{r}||=\nu+\sigma_{r+1}(M(\rho)).$
Thm. 2.2 implies that $\sigma_{r+1}(M(\rho))\leq\nu+\sigma_{r+1}(M)$. Combine the latter two bounds. ∎
@@
Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$, but can be very pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
Remark 2.2.
[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should keep $\rho$ small enough to perform at stage 2 fast or superfast. For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
Remark 2.3.
[A simplified version.] In Sec. 5.3 we test a simplified version of Alg. 2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$. In practice, the choice of integers $\rho$ and then $r$ should adaptive, towards minimizing the error norms of the resulting rank-$\rho$ and rank-$r$ approximations of $M$.
3 LRA by means of random sampling with sparse multipliers
The random sampling LRA algorithms of [11, 33] are close to optimal with a high probability but are not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT. In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations to superfast level by using sparse test matrices, e.g., Ultra-Sparse Rademacher test matrices [33, Sec. 3.9], [3], which depend on a positive integer parameter $s$, such that a test matrices are sparse and support superfast LRA for smaller values $s$. It is a challenge to support our VLRA with these algorithms and test matrices, but in this paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix D and also supporting superfast VLRA. For other candidate sparse test matrices see [11, Remark 4.6], [7, Sec. 5.1.8].
According to [13], such acceleration tends to make the output LRAs somewhat less reliable, and indeed in our tests LRAs computed at stage 1 of Alg. 2.1 tended to be cruder with abridged SRHT test matrices than with Gaussian random ones. Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec. 1.3. We allow at most $k$ refinement steps for a fixed tolerance $k$. As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
@@ 4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
as in Alg. 2.1, except that the input of Alg. 4.1 additionally includes a positive integer $k$, which bounds the number of recursive steps allowed, and an Alg. $\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
@@
    Compute an approximation of $E_{i}$ by SVD of a matrix $E_{i}(\rho)$ of rank $\rho$. [This is step 1 of Alg. 2.1 applied to $E_{i}$ rather than $M$.]
@@
    Compute the matrix $M_{i}(\rho):=X_{i}+E_{i}(\rho)$ by applying Alg. $\mathbb{S}$, performed with a higher precision. [This updates approximation $X_{i}$ to $M$.]
@@
    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This compresses $M_{i}(\rho)$.]
@@
    Output this SVD as an LRA of $M$ and also output the norm $\mu_{i}:=||M-(M_{i}(\rho))_{r}||$ if it satisfies the tolerance bound, that is, if the $\mu_{i}:=\mu_{i}(M,\rho)\leq\mu$.
    Otherwise output FAILURE if $i=k$.
@@
    Otherwise, update the error matrix $E_{i+1}:=M-X_{i+1}$, increase $i$ by 1, and go to stage 1 (to resume the iteration).
@@
[The computational cost bound, the choice of an upper rank, and variations of Alg. 4.1.] As in Alg. 2.1 maximize the upper rank $\rho$ to decrease the output error norm, being only limited by the assumption that the overall computational cost is dominated at stage 1 (cf. (4.1)). We can allow an upper rank $\rho$ to vary as $i$ varies as long that it stays much lower than $n$ for all $i$. In Sec. 5.3 we test a variant of Alg. 4.1 where we fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
In the following extension of Thm. 2.3 we estimate the output error norm of Alg. 4.1.
@@
For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,k-1.$ Then it holds that
$\mu_{i+1}:=||E_{i+1}||\leq\sigma_{r+1}(M)+2\nu_{i}.$ (4.2)
@@
Recall from Alg. 4.1 that
$E_{i+1}:=M-X_{i+1}~{}{\rm for}~{}X_{i+1}:=(X_{i}+E_{i}(\rho))_{r}.$
Hence
$||E_{i+1}||\leq||M-M_{i}(\rho)||+||M_{i}(\rho)-X_{i+1}||=\nu_{i}+\sigma_{r+1}(M_{i}(\rho)).$
Thm. 2.2 implies that
$\sigma_{r+1}(M_{i}(\rho))\leq\sigma_{r+1}(M)+\nu_{i}.$
Combine the latter two bounds. ∎
Remark 4.2.
The upper bound (4.2) on the error norm is within $2\nu_{i}$ from the lower bound $\sigma_{r+1}(M)$, attained where $M_{i}(\rho)=M_{\rho}$.
Next we specify how compression limits the accuracy of the output LRAs.
Lemma 4.1.
For any $m\times n$ matrix $X$ and $m\geq n$ write $\sigma_{j}(X):=\infty$ for $j<1$ and $\sigma_{j}(X):=0$ for $j>n$. Then
[…]
Proof.
Recall that
$E_{i}=M-X_{i}~{}{\rm where}~{}\operatorname{rank}(X_{i})\leq r~{}{\rm for~{}all}~{}i.$
Furthermore,
$\sigma_{j+1}(E_{i})=\min_{Y:~{}\operatorname{rank}(Y)\leq j}||E_{i}-Y||$
by virtue of Thm. 2.1.
Substitute $E_{i}-Y=M-(Y+X_{i})$ where $\operatorname{rank}(Y+X_{i})\leq j+r$ if $\operatorname{rank}(Y)\leq j$ and obtain
$\sigma_{j+1}(E_{i})\geq\min_{Z:~{}\operatorname{rank}(Z)\leq j+r}||M-Z||=\sigma_{j+r+1}(M).$
Similarly deduce that $\sigma_{j+1}(E_{i})\leq\sigma_{j-r+1}(M)$. ∎
@@
The singular values $\sigma_{\rho+1}(E_{i})$ satisfy
@@
[Error norm bounds.] In good accordance with the corollary, in our tests with $(r,\rho)$-matrices the output accuracy of Alg. 2.1 was higher than that of Alg. 4.1 provided that the upper rank $\rho$ in Alg. 2.1 was twice as large as in Alg. 4.1. The test results were mixed for input matrices with flat spectra.